OpenStudy (anonymous):

Solve 24 - 3x = -27.

OpenStudy (anonymous):

24-3x=-27 -24 -24 -3x=-51 / -3 /-3 x=17

OpenStudy (anonymous):

24-3x=-27 -3X=-51 X= 51/3 = 17

OpenStudy (anonymous):

24-3x=-27 24 + 27 = 3x 51 = 3x x = 17